# The Sunburst Arc with JWST. IV. The importance of interaction, turbulence, and feedback for Lyman-continuum escape

### v2

@@ Abstract
At present, the best opportunity for detailed Lyman Continuum escape studies is in gravitationally lensed galaxies at . Only one such galaxy currently exists in the literature with sufficient spatial magnification: The Sunburst Arc at redshift . Here, we present rest-frame optical JWST NIRSpec integral field observations of the Sunburst Arc that cover a large fraction of the source plane. From this dataset, we generate precise maps of ISM kinematics, dust geometry, ionization, and chemical enrichment. We extract a stacked spectrum of five gravitationally lensed images of the Lyman-Continuum leaking cluster, as well as a weighted ,integrated spectrum of most of the galaxy, enabling a direct comparison to other LyC leakers in the literature. We find that the galaxy rotates but also shows strong, possibly dominant, signatures of turbulence, which are indicative of recent or ongoing major interaction. The cluster that leaks ionizing photons shows little variation in kinematics or dust coverage, but dramatically elevated ionization, indicating that photoionization is the predominant mechanism that creates paths for LyC escape. We conjecture that tidal stripping of H I gas due to an interaction could have removed a large portion of the neutral ISM around the LyC emitting cluster, making it easier for the cluster to completely ionize the rest.
@@ 1 Introduction
A few hundred million years after the Big Bang, ionizing photons, termed Lyman-Continuum (LyC) radiation, from the first stars began to ionize the then-neutral IGM (e.g., T. Hashimoto et al., 2018; F. Haardt & P. Madau, 2012; C.-A. Faucher-Giguère, 2020) in what is called the Epoch of Reionization (EoR). Passageways must thus have been present through the neutral ISM of these galaxies, which allowed LyC to leak out into the surrounding IGM, but it is not yet clear how these passageways were carved. In order to have reionized the universe, an estimated fraction of 10–20% of ionizing photons must have escaped their galaxies (B. E. Robertson et al., 2015; R. P. Naidu et al., 2020), with some studies suggesting it might be more (G. D. Becker et al., 2021; F. B. Davies et al., 2021). The ionizing escape fraction cannot be directly measured in the EoR, because the IGM is opaque to ionizing photons at redshifts (A. K. Inoue et al., 2014). All studies of escaping LyC must use lower-redshift galaxies to understand the underlying mechanisms. Such Lyman-Continuum Emitting galaxies (LCE) are rare in the local Universe, but become more common at higher redshifts, as both cosmic star formation activity and ionizing background grow larger. Since the first detection of a Lyman-Continum Emitting galaxy (LCE N. Bergvall et al., 2006), a small number of LCEs have been detected in the local Universe (see e.g. N. Bergvall et al., 2006; E. Leitet et al., 2011; J. Puschnig et al., 2017; G. Östlin et al., 2021; L. Komarova et al., 2024; C. Leitherer et al., 2016); and at slightly higher redshifts, a considerable number of Green Pea C. Cardamone et al. (2009) and similar galaxies have found to be LCEs (e.g. S. Borthakur et al., 2014; A. E. Jaskot et al., 2019; B. Wang et al., 2021; M. A. Malkan & B. K. Malkan, 2021; S. R. Flury et al., 2022a). At redshifts , about LCEs have been identified, depending on delineation criteria for candidates (E. Vanzella et al., 2012; R. E. Mostardi et al., 2015; S. de Barros et al., 2016; A. E. Shapley et al., 2016; E. Vanzella et al., 2016; F. Bian et al., 2017; J. Chisholm et al., 2018; T. J. Fletcher et al., 2019; C. C. Steidel et al., 2018; E. Vanzella et al., 2018; T. E. Rivera-Thorsen et al., 2019, 2022; K. Saha et al., 2020; R. Marques-Chaves et al., 2021, 2022; Z. Ji et al., 2020; S. R. Flury et al., 2022a; A. J. Pahl et al., 2021; A. Saxena et al., 2022; J. Kerutt et al., 2024).
@@
Since it is not possible to directly measure galaxy LyC emission during the EoR, much work has been put into the identification and characterization of secondary observables which correlate with LyC escape, and the dual, intertwined questions of a) how they statistically scale with , and b) what they reveal about the mechanisms that carve out the paths of escape for the ionizing radiation. Very simplified, these mechanisms fall into the families of radiative versus mechanical feedback, which give rise to different escape scenarios. In a spherically symmetric, dust-free toy model of a young star cluster enveloped in an isotropic, neutral ISM, photoionization over time leads to a falling neutral fraction of the ISM in a growing Strömgren sphere which may eventually grow larger than the boundary of the ISM. This scenario is often called the “ionization bounded” scenario, where there is still an intact neutral envelope present, and the “density bounded scenario” when the Strömgren sphere has grown larger than the gas envelope (see e.g. E. Zackrisson et al., 2013, for an overview). In such a scenario, depends solely on the column density of the residual neutral gas fraction in the ISM.
@@
In contrast, radiative pressure and mechanical feedback such as stellar winds or momentum deposited from Supernovae, can result in large-scale ISM outflows, in which Rayleigh-Taylor instabilities will result in a breakup of this expanding bubble into a clumpy, highly anisotropic, expanding medium, in which some lines of sight will be transparent to ionizing photons, while others are optically thick. This is often referred to as the “perforated ionization bounded scenario” or the “picket fence scenario” (e.g. T. M. Heckman et al., 2011; E. Zackrisson et al., 2013; B. L. James et al., 2014; A. E. Jaskot & M. S. Oey, 2014); in this limit, is regulated solely by the global or line-of-sight (LOS) covering fraction of the neutral gas. Real-life scenarios are of course much more complex, but the escape fraction can still be considered as regulated by the covering fraction of neutral gas clumps and the residual neutral column in the inter-clump medium.
@@
A number of observables either correlate with, or directly encode, these two properties. One of the most direct measurements of the line-of-sight covering fraction is the equivalent width of H I absorption lines (T. M. Heckman et al., 2011; T. A. Jones et al., 2013; S. Gazagnes et al., 2018; J. Chisholm et al., 2018), although for practical reasons, neutral metal lines are often used as proxies for these (T. E. Rivera-Thorsen et al., 2017a; R. M. Alexandroff et al., 2015; T. A. Jones et al., 2013; S. Borthakur et al., 2014; T. M. Heckman et al., 2011, e.g.). However, absorption lines are very observationally expensive, especially at high redshifts and thus often not practical for EoR studies. While neutral absorption lines directly trace the covering fraction, the properties of the Lyman- line directly traces the neutral column density in the inter-clump medium, although the strongly resonant nature of the line allows it to traverse open channels without any direct open lines of sight (M. Gronke et al., 2016; M. Dijkstra et al., 2016; M. Dijkstra, 2014; A. Verhamme et al., 2015, 2008, 2006, e.g.). A number of line properties have been found to correlate with LyC escape, such as e.g. velocity offset of the red peak or separation of the dual peaks (A. Verhamme et al., 2008, 2015; A. E. Jaskot et al., 2024a, b), the steepness of the red peak (K. Kakiichi & M. Gronke, 2021). The presence of a narrow line component at systemic velocity has been predicted (C. Behrens et al., 2014; A. Verhamme et al., 2015; T. E. Rivera-Thorsen et al., 2017a) and confirmed (T. E. Rivera-Thorsen et al., 2019) to indicate the presence of a line of sight of extremely low . However, due to the predominantly neutral IGM at the EoR, Ly emission cannot be used to trace ionizing escape at that time.
@@
Other observables correlate either with high ionizing production rate, high escape fractions, or with physical conditions that bring about these conditions. In the latter category is the finding that LyC escape seems to correlate with broad emission line components, which in turn indicate the presence of large scale bulk outflows which can rupture and thin out the neutral ISM and create escape channels (see e.g. R. Mainali et al., 2022, and references therein). However, in a study of low-redshift extreme emission line galaxies (EELGs), A. E. Jaskot et al. (2019) found that outflows tended to be weaker at higher ionization, which they interpret as the onset of catastrophic cooling of neutral gas clumps which subsequently are photoionized by the still-proximate ionizing sources; marking the transition to a regime dominated by the density bounded scenario. LyC escape has also been found to correlate with star formation surface density (e.g. A. E. Jaskot et al., 2024a, and references therein).
@@
Besides properties internal to the source galaxy, the impact of the environment is a topic of active debate and research. Major mergers and IGM accretion events are predicted to spark strong star formation episodes, which in turn would increase the ionizing photon production, and feedback from this star formation could lead to increased escape fractions (e.g. I. Kostyuk & B. Ciardi, 2024). Major interactions could also lead to strong turbulence, which in simulations have shown to be able to create ionizing escape paths through the ISM (e.g., K. Kakiichi & M. Gronke, 2019).
@@
Observations have been ambiguous as to the importance of mergers. S. Zhu et al. (2025) find from visual morphological classification of a sample of 23 LCEs in GOODS-South that 20 of them are mergers, a dramatic overrepresentation relative to the general merger fraction at (Q. Duan et al., 2025; D. Puskás et al., 2025). In their simulated high-redshift IllustrisTNG galaxies, I. Kostyuk & B. Ciardi (2024) find that not only star formation induced by mergers, but also tidal stripping of the surrounding H I envelope plays a crucial role in facilitating LyC escape. Interestingly, this effect was observed directly by A. Le Reste et al. (2024), who used 21 cm interferometry to directly map H I in the local-Universe, interacting LCE Haro 11. These authors found that tidal stripping by the companion galaxy had offset large portions of the H I envelope away from the star forming knots in that galaxy, helping to clear out the ionizing escape paths.
@@
At higher redshifts, where LyC has been shifted into the detectable wavelength ranges of HST imaging instruments, spatial detail is lost to distance. However, magnification by gravitational lensing can in fortuitous cases image galaxies at cosmological distances at a level of detail that is usually only seen in the local Universe. At the same time, at these distances, the ionizing radiation has been redshifted into ranges where it is comfortably observable in the UVIS channel of HST, at physical de-magnified resolutions down to a few tens of parsecs or occasionally even better. As a result, such gravitationally lensed LCE galaxies currently present the best available opportunity for this kind of high-detail mapping of ionizing escape paths.
@@
However, such lensed LCEs are rare. While searches are ongoing, at the time of writing, only one lensed LCE is found in the literature, the Sunburst Arc (catalog ) (H. Dahle et al., 2016; T. E. Rivera-Thorsen et al., 2019) at redshift . The Sunburst Arc was discovered as part of the Planck foreground project by H. Dahle et al. (2016), who found it to be the brightest known gravitationally lensed galaxy with an integrated magnitude of . Lensing by a redshift foreground cluster generates a total of 12 total or partial images (H. Dahle et al., 2016; K. Sharon et al., 2022; G. V. Pignataro et al., 2021; J. M. Diego et al., 2022; E. Solhaug et al., 2025). T. E. Rivera-Thorsen et al. (2017b) found that the galaxy displayed a unique, triple-peaked Ly emission profile, which is predicted to arise from Ly radiative transfer in an expanding shell of Hi perforated by a narrow channel devoid of Hi along our line of sight (C. Behrens et al., 2014; A. Verhamme et al., 2015). T. E. Rivera-Thorsen et al. (2017b) estimated the neutral column density along this line of sight to be extremely low, , while one optical depth for LyC corresponds to ; orders of magnitude higher. The galaxy is dominated by a very young (3–5 Myr J. Chisholm et al., 2019; E. Vanzella et al., 2020; M. Pascale et al., 2023; U. Meštrić et al., 2023; T. E. Rivera-Thorsen et al., 2024) and massive ( E. Vanzella et al., 2020; T. E. Rivera-Thorsen et al., 2024) Lyman Continuum emitting (T. E. Rivera-Thorsen et al., 2019, from here referred to as "the LCE") star cluster. This cluster is dense and massive and contains a large population of Wolf-Rayet stars and Very Massive Stars (M. Pascale et al., 2023; U. Meštrić et al., 2023; T. E. Rivera-Thorsen et al., 2024), and a strong enrichment in both doubly (M. Pascale et al., 2023) and singly (B. Welch et al., 2025) ionized Nitrogen, as has also been observed in a number of high-redshift galaxies with JWST (e.g. Z. Ji et al., 2020, and references therein). The LCE cluster has a very steep rest-frame UV slope of (K. J. Kim et al., 2023), and R. Mainali et al. (2022) found strong broad emission line components in ground-based, rest-frame optical spectroscopy, which those authors interpreted as the presence of strong bulk outflows driven by stellar winds from the young stellar population of the cluster. Based on this, R. Mainali et al. concluded that mechanical feedback was the main driver of ionizing escape in the Sunburst Arc, presumably by driving the expansion and rupturing the surrounding neutral ISM. The present NIRSpec IFU observations provide an opportunity to spatially map the kinematic properties which have previously only been attainable in integrated slit spectra through atmospheric distortion, among other things enabling us to test this scenario more thoroughly.
@@
This paper assumes a standard -CDM cosmology with km/s/Mpc, . Unless otherwise noted, all images are oriented with N up and E left. We use the abbreviations [O I] for [O I] 6300, [O III] for [O III] 5008, [O II] for [O II] 3727+29, [N II] for [N II] 6584, [Ne III] for [Ne III] 3869, [S II] for [S II] 6717+31.
@@ 2 Observations
The JWST (J. P. Gardner et al., 2023) data analyzed in this paper have also been included in the previous papers in this series; T. E. Rivera-Thorsen et al. (2024) , (Paper I), S. Choe et al. (2025)(Paper II), and B. Welch et al. (2025). The observations were obtained with the NIRSpec instrument (T. Böker et al., 2023) in the IFS mode with the G140H and G235H gratings, and and with the NIRCam instrument (M. J. Rieke et al., 2023). These observations were obtained over the period from April 10, 2023, to April 4, 2023, as part of program GO-2555 (PI Rivera-Thorsen).
@@
The NIRSpec pointings were designed such that P1 covers the largest possible part of the galaxy in one pointing; P2 is centered around the peculiar clump in its center, denoted “Tr” in E. Vanzella et al. (2020), or “Godzilla” by J. M. Diego et al. (2022); and P3 is placed to cover two strongly magnified images of the LCE cluster, to optimize the amount of information we can get about this cluster and its closest surroundings. Since the lensing shear of P1 is approximately perpendicular to the lensing shear in P2 and P3 (according to the lensing model of K. Sharon et al. 2022), this choice of pointings optimizes the amount of spatial information that can be recovered.
Fig. 2 shows an artistic approximation of the de-lensed Sunburst Galaxy based on HST imaging, identical to the left panel of Fig. 11 in K. Sharon et al. (2022), except here we have overlaid the approximate de-lensed IFU footprints; the direction in which each distorted square is shortest is the direction of its maximal spatial resolution. The figure consists of an image of the full, de-lensed arc as seen in the W and SE arc segments in K. Sharon et al. (see 2022, Fig. 2), onto which are superimposed by hand the more precise positions, colors, and brightnesses of the better resolved clumps found in the N and NW arc segments. The redder part in the N is a likely interacting companion galaxy, while the blue-green blurry parts in the center are likely to be a tidal bridge or similar feature.
NIRSpec Integral Field Unit (IFU) data were reduced following the methods presented in J. R. Rigby et al. (2023). We used the JWST data reduction pipeline version 1.11.4 (H. Bushouse et al., 2023), with the calibration reference files from context pmap_1105. As discussed in J. R. Rigby et al. (2023), the electronics of the NIRSpec detectors undergo small-scale fluctuations in temperature, which create pattern noise. We correct this detector pattern noise using the NSClean software (B. J. Rauscher, 2023).
The raw NIRSpec data contains a large number of outlier pixels due to cosmic rays. The standard pipeline includes an outlier rejection step; however, this step does not adequately remove outliers from our final data products. We therefore use the baryon-sweep code (T. A. Hutchison et al., 2024) to remove remaining outliers and artifacts. T. A. Hutchison et al. (2023) found that a combination of the standard pipeline outlier rejection and the baryon-sweep method produces the cleanest final data products.
@@
Single-spaxel spectra of point-like sources from either of JWST’s integral field units (MIRI MRS and the NIRspec IFU) are susceptible to a “wiggle” artifact showing as faint, very broad waves in the continuum , which arises in each case because the detector does not critically sample the point spread function (D. R. Law et al., 2023). For the NIRSpec IFU, drizzling this undersampled data to even smaller output spaxels exaggerates the artifact (when using a standard 4-point dither as was done in this program). As noted by D. R. Law et al. (2023), using a larger extraction aperture around point sources mitigates this artifact, even in the NIRSpec data cubes. We find that the benefit from improved spatial resolution for extended sources outweighs the additional artifacts seen in point-like sources within this dataset, have opted to analyze the 0.″05 data cubes.

### v3

@@ Abstract
At present, the best opportunity for detailed Lyman Continuum escape studies is in gravitationally lensed galaxies at . Only one such galaxy currently exists in the literature with sufficient magnification and spatial resolution: The Sunburst Arc at redshift . Here, we present rest-frame optical JWST NIRSpec integral field observations of the Sunburst Arc that cover a large fraction of the source plane. From this dataset, we generate precise maps of ISM kinematics, dust geometry, ionization, and chemical enrichment. We extract a stacked spectrum of five gravitationally lensed images of the Lyman-Continuum leaking star cluster, as well as a weighted, integrated spectrum of most of the galaxy, enabling a direct comparison to other LyC leakers in the literature. We find that the galaxy rotates but also shows strong, possibly dominant, signatures of turbulence, which are indicative of recent or ongoing major interaction. The cluster that leaks ionizing photons shows little difference from the rest of the galaxy in terms of kinematics or dust coverage, but dramatically elevated ionization, indicating that photoionization is the predominant mechanism that creates paths for LyC escape. We propose that tidal stripping of H I gas due to an interaction may have removed a large portion of the neutral ISM around the LyC emitting cluster, making it easier for the cluster to completely ionize the rest.
@@ 1 Introduction
A few hundred million years after the Big Bang, ionizing photons, termed Lyman-Continuum (LyC) radiation, from the first stars began to ionize the then-neutral IGM (e.g., hashimoto2018nature; haardt2012; fauchergiguere2020) in what is called the Epoch of Reionization (EoR). Passageways must thus have been present through the neutral ISM of these galaxies, which allowed LyC to leak out into the surrounding IGM, but it is not yet clear how these passageways were carved. In order to have reionized the universe, an estimated fraction of 10–20% of ionizing photons must have escaped their galaxies (robertson2015; naidu2020a), with some studies suggesting it might be more (becker2021; davies2021). The ionizing escape fraction cannot be directly measured in the EoR, because the IGM is opaque to ionizing photons at redshifts (inoue2014). All studies of escaping LyC must use lower-redshift galaxies to understand the underlying mechanisms. Such Lyman-Continuum Emitting galaxies (LCE) are rare in the local Universe, but become more common at higher redshifts, as both cosmic star formation activity and ionizing background grow larger. Since the first detection of a Lyman-Continum Emitting galaxy (LCE; bergvall2006), a small number of LCEs have been detected in the local Universe (see e.g. bergvall2006; leitet2011; puschnig2017; ostlin2021; komarova2024; leitherer2016); and at slightly higher redshifts, a considerable number of Green Pea (cardamone2009) and similar galaxies have found to be LCEs (e.g. borthakur2014; jaskot2019; wang2021; malkan2021b; flury2022a). At redshifts , about LCEs have been identified, depending on delineation criteria for candidates (vanzella2012; mostardi2015; debarros2016; shapley2016; vanzella2016; bian2017; chisholm2018a; fletcher2018; steidel2018; vanzella2018; riverathorsen2019; riverathorsen2022; saha2020; marqueschaves2021; marqueschaves2022; ji2020; flury2022a; pahl2021; saxena2022; kerutt2024).
@@
Since it is not possible to directly measure galaxy LyC emission during the EoR, much work has been put into the identification and characterization of secondary observables which correlate with LyC escape, and the dual, intertwined questions of a) how they statistically scale with , and b) what they reveal about the mechanisms that carve out the paths of escape for the ionizing radiation. Very simplified, these mechanisms fall into the families of radiative versus mechanical feedback, which give rise to different escape scenarios. In a spherically symmetric, dust-free toy model of a young star cluster enveloped in an isotropic, neutral ISM, photoionization over time leads to a falling neutral fraction of the ISM in a growing Strömgren sphere which may eventually grow larger than the boundary of the ISM. This scenario is often called the “ionization bounded” scenario, where there is still an intact neutral envelope present, and the “density bounded scenario” when the Strömgren sphere has grown larger than the gas envelope (see e.g. zackrisson2013, for an overview). In such a scenario, depends solely on the column density of the residual neutral gas fraction in the ISM.
@@
In contrast, radiative pressure and mechanical feedback such as stellar winds or momentum deposited from Supernovae, can result in large-scale ISM outflows, in which Rayleigh-Taylor instabilities will result in a breakup of this expanding bubble into a clumpy, highly anisotropic, expanding medium, in which some lines of sight will be transparent to ionizing photons, while others are optically thick. This is often referred to as the “perforated ionization bounded scenario” or the “picket fence scenario” (e.g. heckman2011; zackrisson2013; james2014; jaskot2014); in this limit, is regulated solely by the global or line-of-sight (LOS) covering fraction of the neutral gas. Real-life scenarios are of course much more complex, but the escape fraction can still be considered as regulated by the covering fraction of neutral gas clumps and the residual neutral column in the inter-clump medium.
@@
A number of observables either correlate with, or directly encode, these two properties. One of the most direct measurements of the line-of-sight covering fraction is the equivalent width of H I absorption lines (heckman2011; jones2013; gazagnes2018; chisholm2018a), although for practical reasons, neutral metal lines are often used as proxies for these (e.g. riverathorsen2017; alexandroff2015; jones2013; borthakur2014; heckman2011). However, absorption lines are very observationally expensive, especially at high redshifts and thus often not practical for EoR studies. While neutral absorption lines directly trace the covering fraction, the properties of the Lyman- line directly traces the neutral column density in the inter-clump medium, although the strongly resonant nature of the line allows it to traverse open channels without any direct open lines of sight (e.g. gronke2016letter; dijkstra2016; dijkstrarev; verhamme2015; verhamme2008; verhamme2006). A number of line properties have been found to correlate with LyC escape, such as e.g. velocity offset of the red peak or separation of the dual peaks (verhamme2008; verhamme2015; jaskot2024a; jaskot2024b), or the steepness of the red peak (kakiichi2021). The presence of a narrow line component at systemic velocity has been predicted (behrens2014; verhamme2015; riverathorsen2017) and confirmed (riverathorsen2019) to indicate the presence of a line of sight of extremely low . However, due to the predominantly neutral IGM at the EoR, Ly emission cannot be depended on to trace ionizing escape at that time.
@@
Other observables correlate either with high ionizing production rate, high escape fractions, or with physical conditions that bring about these conditions. In the latter category is the finding that LyC escape seems to correlate with broad emission line components, which in turn indicate the presence of large scale bulk outflows which can rupture and thin out the neutral ISM and create escape channels (see e.g. mainali2022, and references therein). However, in a study of low-redshift extreme emission line galaxies (EELGs), jaskot2019 found that outflows tended to be weaker at higher ionization, which they interpret as the onset of catastrophic cooling of neutral gas clumps which subsequently are photoionized by the still-proximate ionizing sources; marking the transition to a regime dominated by the density bounded scenario. LyC escape has also been found to correlate with star formation surface density (e.g. jaskot2024a, and references therein).
@@
Besides properties internal to the source galaxy, the impact of the environment is a topic of active debate and research. Major mergers and IGM accretion events are predicted to spark strong star formation episodes, which in turn would increase the ionizing photon production, and feedback from this star formation could lead to increased escape fractions (e.g. kostyuk2024). Major interactions could also lead to strong turbulence, which in simulations have shown to be able to create ionizing escape paths through the ISM (e.g., kakiichi2019).
@@
Observations have been ambiguous as to the importance of mergers. zhu2025 find from visual morphological classification of a sample of 23 LCEs in GOODS-South that 20 of them are mergers, a dramatic overrepresentation relative to the general merger fraction at (duan2025; puskas2025). In their simulated high-redshift IllustrisTNG galaxies, kostyuk2024 find that not only star formation induced by mergers, but also tidal stripping of the surrounding H I envelope plays a crucial role in facilitating LyC escape. Interestingly, this effect was observed directly by lereste2024, who used 21 cm interferometry to directly map H I in the local-Universe, interacting LCE Haro 11. These authors found that tidal stripping by the companion galaxy had offset large portions of the H I envelope away from the star forming knots in that galaxy, helping to clear out the ionizing escape paths.
@@
At higher redshifts, where LyC has been shifted into the detectable wavelength ranges of HST imaging instruments, spatial detail is lost to distance. However, magnification by gravitational lensing can in fortuitous cases image galaxies at cosmological distances at a level of detail that is usually only seen in the local Universe. At the same time, at these distances, the ionizing radiation has been redshifted into ranges where it is comfortably observable in the WFC3 UVIS channel of HST, at physical de-magnified resolutions down to a few tens of parsecs or occasionally even better. As a result, such gravitationally lensed LCE galaxies currently present the best available opportunity for this kind of high-detail mapping of ionizing escape paths.
@@
However, such lensed LCEs are rare. While searches are ongoing, at the time of writing, only one lensed LCE of sufficient brightness for detailed spatially and spectrally resolved studies is found in the literature; the Sunburst Arc (catalog ) (dahle2016; riverathorsen2019) at redshift . The Sunburst Arc was discovered during optical follow-up of Sunyaev-Zeldovich galaxy cluster candidates in the ESA Planck Survey (plancksz2014) by dahle2016, who found it to be the brightest known gravitationally lensed galaxy with an integrated magnitude of . Lensing by a redshift foreground cluster generates a total of 12 total or partial images (dahle2016; sharon2022; pignataro2021; diego2022; solhaug2025). sunburst2017 found that the galaxy displayed a unique, triple-peaked Ly emission profile, which is predicted to arise from Ly radiative transfer in an expanding shell of Hi perforated by a narrow channel devoid of Hi along our line of sight (behrens2014; verhamme2015). sunburst2017 estimated the neutral column density along this line of sight to be extremely low, , while one optical depth for LyC corresponds to ; orders of magnitude higher. The galaxy is dominated by a very young (3–5 Myr chisholm2019; vanzella2020; pascale2023; mestric2023; riverathorsen2024) and massive ( vanzella2020; riverathorsen2024) Lyman Continuum emitting (riverathorsen2019, from here referred to as "the LCE") star cluster. This cluster is dense and massive and contains a large population of Wolf-Rayet stars and Very Massive Stars (pascale2023; mestric2023; riverathorsen2024), and a strong enrichment in both doubly (pascale2023) and singly (welch2025) ionized Nitrogen, as has also been observed in a number of high-redshift galaxies with JWST (e.g. ji2025arxiv, and references therein). The LCE cluster has a very steep rest-frame UV slope of (kim2023), and mainali2022 found strong broad emission line components in ground-based, rest-frame optical spectroscopy, which those authors interpreted as the presence of strong bulk outflows driven by stellar winds from the young stellar population of the cluster. Based on this, mainali2022 concluded that mechanical feedback was the main driver of ionizing escape in the Sunburst Arc, presumably by driving the expansion and rupturing the surrounding neutral ISM. The present NIRSpec IFU observations provide an opportunity to spatially map the kinematic properties which have previously only been attainable in integrated slit spectra through atmospheric distortion, among other things enabling us to test this scenario more thoroughly.
@@
This paper assumes a standard -CDM cosmology with km s-1 Mpc-1, . Unless otherwise noted, all images are oriented with N up and E left. We use the abbreviations [O I] for [O I] 6300, [O III] for [O III] 5008, [O II] for [O II] 3727+29, [N II] for [N II] 6584, [Ne III] for [Ne III] 3869, [S II] for [S II] 6717+31.
@@ 2 Observations
The JWST (gardner2023) data analyzed in this paper have also been included in the previous papers in this series; riverathorsen2024 , (riverathorsen2024), choe2025(choe2025), and welch2025. The observations were obtained with the NIRSpec instrument (boeker2023) in the IFS mode with the G140H and G235H gratings, and and with the NIRCam instrument (rieke2023). These observations were obtained over the period from April 10, 2023, to April 4, 2023, as part of program GO-2555 (PI Rivera-Thorsen).
@@
The NIRSpec pointings were designed such that P1 covers the largest possible part of the galaxy in one pointing; P2 is centered around the peculiar clump in its center, denoted “Tr” in vanzella2020, or “Godzilla” by diego2022; and P3 is placed to cover two strongly magnified images of the LCE cluster, to optimize the amount of information we can get about this cluster and its closest surroundings. Since the lensing shear of P1 is approximately perpendicular to the lensing shear in P2 and P3 (according to the lensing model of sharon2022), this choice of pointings optimizes the amount of spatial information that can be recovered.
Fig. 2 shows an artistic approximation of the de-lensed Sunburst Galaxy based on HST imaging, identical to the left panel of Fig. 11 in sharon2022, except here we have overlaid the approximate de-lensed IFU footprints; the direction in which each distorted square is shortest is the direction of its maximal spatial resolution. The figure consists of an image of the full, de-lensed arc as seen in the W and SE arc segments in sharon2022, onto which are superimposed by hand the more precise positions, colors, and brightnesses of the better resolved clumps found in the N and NW arc segments. The redder part in the N is a likely interacting companion galaxy, while the blue-green blurry parts in the center are likely to be a tidal bridge or similar feature.
NIRSpec Integral Field Unit (IFU) data were reduced following the methods presented in rigby23_overview. We used the JWST data reduction pipeline version 1.11.4 (bushouse23_pipeine1p11p4), with the calibration reference files from context pmap_1105. As discussed in rigby23_overview, the electronics of the NIRSpec detectors undergo small-scale fluctuations in temperature, which create pattern noise. We correct this detector pattern noise using the NSClean software (rauscher23_nsclean).
The raw NIRSpec data contains a large number of outlier pixels due to cosmic rays. The standard pipeline includes an outlier rejection step; however, this step does not adequately remove outliers from our final data products. We therefore use the baryon-sweep code (baryonsweep) to remove remaining outliers and artifacts. hutchison23_sigmaclip found that a combination of the standard pipeline outlier rejection and the baryon-sweep method produces the cleanest final data products.
@@
Single-spaxel spectra of point-like sources from either of JWST’s integral field units (MIRI MRS and the NIRspec IFU) are susceptible to a “wiggle” artifact showing as faint, very broad waves in the continuum , which arises in each case because the detector does not critically sample the point spread function (law23_3ddrizzle). These wiggles are much broader than emission lines and easily removed in continuum subtraction, but they do give rise to these Moiré-like patterns in ratio maps. For the NIRSpec IFU, drizzling this undersampled data to even smaller output spaxels exaggerates the artifact (when using a standard 4-point dither as was done in this program). As noted by law23_3ddrizzle, using a larger extraction aperture around point sources mitigates this artifact, even in the NIRSpec data cubes. We find that the benefit from improved spatial resolution for extended sources outweighs the additional artifacts seen in point-like sources within this dataset. Hence, we have opted to analyze the 0.″05 data cubes.